Trish's veterinarian diagnosed her cat with feline acne. He is out of the 1% benzoyl peroxide needed to treat it. So he recommended that she buy the 5% benzoyl peroxide for human acne in the drug store and dilute it to 1%. How many teaspoons of water should Trish mix with 1 teaspoon of the 5% benzoyl peroxide to get 1% benzoyl peroxide?
Answer by stanbon(75887)

He is out of the 1% benzoyl peroxide needed to treat it. So he recommended that she buy the 5% benzoyl peroxide for human acne in the drug store and dilute it to 1%. How many teaspoons of water should Trish mix with 1 teaspoon of the 5% benzoyl peroxide to get 1% benzoyl peroxide?
--------------------
Equation:
benz + benz = benz
0.05*1 + 0*x = 0.01(1+x)
----
Multiply thru by 100 to get:
5*1 + 0 = 1*1 + x
x = 4 (# of teaspoons of Water needed)
------
Cheers,
Stan H.
-----------------